QN= 100-0.25p
Qw=200-0.5p
The school will admit 180 students from each group (students may choose not to enroll). The marginal cost to the school is constant at Shs 1000 per student course hour.UON'S fixed cost is Shs 2.0million per year.
For each of the following pricing schemes, please derive the optimal price structure and the total profit on the assumption that the UON's objective is to Maximize profit for the program.A single tuition per academic year plus a single fee per course hour.
1
Expert's answer
2020-04-20T10:29:38-0400

If QN = 100 - 0.25p, then p = 400 - 4QN.

If Qw = 200 - 0.5p, then p = 400 - 2Qw.

The profits are maximized when MR = MC.

"MRN = TR'(QN) = 400 - 8QN,"

"MRw = TR'(Qw) = 400 - 4Qw,"

"400 - 8QN = 1000,"

QN is negative.

400 - 4Qw = 1000,

Qw is also negative, so nobody will enroll.